OpenStudy (anonymous):

suppose you have a bag that holds 30 pieces of candy: 3 blueberry, 19 cream soda, and 8 apple. you reach into the bag to pull out one piece. find the probability that the piece you pull out is not apple flavored.

OpenStudy (anonymous):

P(no apple)=1-P(apple)

OpenStudy (callisto):

P(not apple) = (blueberry + cream soda)/ total = (3+19) / 30 = 22/30 = 11/15

OpenStudy (anonymous):

there are 22 out of 30 that arent apple 22/30 = 11/15

hero (hero):

Good job @Callisto

OpenStudy (callisto):

Nah?! I think I was kinda stupid not to use 1- P(apple) :(. But thanks!

hero (hero):

Why do you say that?

OpenStudy (callisto):

Somehow, thinking in another way is faster

hero (hero):

I think your way is perfectly fine. In this case, a clear answer supercedes the fast answer.

OpenStudy (callisto):

Alright , thanks!